# The First Constraint on the Dark Matter Millicharge via Multimessenger Time-Delay Analysis of GRB GW170817A

### v1

@@ I Introduction
Millicharged particles, which carry a fraction of electron charge, $Q=\epsilon e$, are proves of the violation of charge-quantization hypothesis and predicted by the superstring theories [1, 2] and grand unified theories [3, 4]. The dark sector provides a natural scenario for weakly breaking the U(1) electromagnetic symmetry through interactions with hidden photon fields [5, 6, 7]. The interaction between millicharged Dark Matter (DM) and the Standard Model (SM) is mediated by a kinetically mixed dark photon [8, 9, 10], effectively introducing a feeble U(1) coupling. For wave-like dark matter with mass below $10^{-10}$ eV, terrestrial detectors such as the milliQan experiment [11], colliders [12, 13, 14], fixed-target experiments [15, 16, 17, 18, 17, 19] and neutrino experiments [20, 21, 22], face challenges due to the ultralight particle’s macroscopic de Broglie wavelength [23], necessitating probes via spectral distortions in astrophysical or cosmological transient events [13]. In this paper, we do not address the production mechanisms of millicharged dark matter, but instead propose a novel detection scheme for light dark matter in the Universe. This approach leverages the observed time delay between the short gamma-ray burst (sGRB) and gravitational wave emission from the neutron star merger event [24, 25, 26], GW170817 [27, 28, 29]. By treating the ubiquitous millicharged dark matter in the cosmos as a homogeneous medium — analogous to light propagation in water where photon scattering with the medium slows the effective phase velocity — we analyze how interactions between photons and millicharged dark matter could induce a measurable delay in high-energy photon arrival times.
@@ II Light velocity in DM medium and Constraints from the sGRB GW170817A
  • Scenario2: Assuming the time lag $\sim 0.01$s between (8, 50) keV and (100, 300) keV photons arriving at Earth is induced by scattering with dark matter. In this case, the sGRB and gravitational waves are not produced simultaneously - the sGRB emission lags behind the GWs by approximately one second, which is consistent with predictions from jet formation models for sGRBs. The time lag induced by the millichared DM is no longer than $\sim 0.27$s. Approximately, adopting $30$ keV and $200$ keV as representative energies for (8, 50) keV and (100, 300) keV bands respectively, we derive
@@ V Appendix
Here, we omit the field operators unrelated to the creation or annihilation of ALPs. Field operators connected by contraction lines must undergo contraction, while those without contraction symbols participate in the creation and annihilation of initial and final state particles. In the second line, the factor, $N_{i}^{2}$, arises from the number of momentum eigenstates p and q selected from the initial and final states. The last line in the eq.(V.5) is obtained by using $<q_{k}|p_{j}>=(2\pi)^{3}(2E_{p_{j}})\delta^{3}(q_{k}-p_{j})$. After integrating over the momentum variables $q_{k}$, this term acquires an $N_{i}-1!$ factor stemming from the permutation symmetry of the BEC particles’ momenta between the initial and final states. Therefore,

### v2

@@ I Introduction
Millicharged particles, which carry a fraction of electron charge, $Q=\epsilon e$, are proves of the violation of charge-quantization hypothesis and predicted by the superstring theories [1, 2] and grand unified theories [3, 4]. The dark sector provides a natural scenario for weakly breaking the U(1) electromagnetic symmetry through interactions with hidden photon fields [5, 6, 7]. The interaction between millicharged Dark Matter (DM) and the Standard Model (SM) is mediated by a kinetically mixed dark photon [8, 9, Fabbrichesi:2020wbt,Arza:2025cou], effectively introducing a feeble U(1) coupling. For wave-like dark matter with mass below $10^{-10}$ eV, terrestrial detectors such as the milliQan experiment [10], colliders [11, 12, 13], fixed-target experiments [14, 15, 16, 17, 16, 18] and neutrino experiments [19, 20, 21], face challenges due to the ultralight particle’s macroscopic de Broglie wavelength [22], necessitating probes via spectral distortions in astrophysical or cosmological transient events [12]. In this paper, we do not address the production mechanisms of millicharged dark matter, but instead propose a novel detection scheme for light dark matter in the Universe. This approach leverages the observed time delay between the short gamma-ray burst (sGRB) and gravitational wave emission from the neutron star merger event [23, 24, 25], GW170817 [26, 27, 28]. By treating the ubiquitous millicharged dark matter in the cosmos as a homogeneous medium — analogous to light propagation in water where photon scattering with the medium slows the effective phase velocity — we analyze how interactions between photons and millicharged dark matter could induce a measurable delay in high-energy photon arrival times.
@@ II Light velocity in DM medium and Constraints from the sGRB GW170817A
  • Scenario2: Assuming the time lag $\sim 0.27$s between (8, 50) keV and (100, 300) keV photons arriving at Earth is induced by scattering with dark matter. In this case, the sGRB and gravitational waves are not produced simultaneously - the sGRB emission lags behind the GWs by approximately one second, which is consistent with predictions from jet formation models for sGRBs. The time lag induced by the millichared DM is no longer than $\sim 0.27$s. Approximately, adopting $30$ keV and $200$ keV as representative energies for (8, 50) keV and (100, 300) keV bands respectively, we derive
@@ V Appendix
Here, we omit the field operators unrelated to the creation or annihilation of ALPs. Field operators connected by contraction lines must undergo contraction, while those without contraction symbols participate in the creation and annihilation of initial and final state particles. In the second line, the factor, $N_{i}^{2}$, arises from the number of momentum eigenstates p and q selected from the initial and final states. The last line in the eq.(V.5) is obtained by using $<q_{k}|p_{j}>=(2\pi)^{3}(2E_{p_{j}})\delta^{3}(q_{k}-p_{j})$. After integrating over the momentum variables $q_{k}$, this term acquires an $N_{i}-1!$ factor stemming from the permutation symmetry of the BEC particles’ momenta between the initial and final states and finally leaves an enhancement factor of $N_{i}$.